OpenStudy (anonymous):

The measure of a side of a square field is 2 x squared + 3. What is its perimeter?

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (skullpatrol):

What does the word "perimeter" mean?

OpenStudy (anonymous):

the continuous line forming the boundary of a closed geometric figure

OpenStudy (skullpatrol):

Yes, more specifically it is the length of that line. perimeter: The perimeter of a plane figure is the distance around it.

OpenStudy (anonymous):

so 2^2 ( 2times 2) = 4x^2 .. ( 3^2) =9 ....4x^2 = 9 ?

OpenStudy (anonymous):

4x^2 + 9 ?

OpenStudy (skullpatrol):